# Evaluating machine learning models for predicting pesticides toxicity to honey bees

### v3

@@ Abstract
Small molecules play a critical role in the biomedical, environmental, and agrochemical domains, each with distinct physicochemical requirements and success criteria. Although biomedical research benefits from extensive datasets and established benchmarks, agrochemical data remain scarce, particularly with respect to species-specific toxicity. This work focuses on ApisTox, the most comprehensive dataset of experimentally validated chemical toxicity to the honey bee (Apis mellifera), an ecologically vital pollinator. We evaluate ApisTox using a diverse suite of machine learning approaches, including molecular fingerprints, graph kernels, and graph neural networks, as well as pretrained models. Comparative analysis with medicinal datasets from the MoleculeNet benchmark reveals that ApisTox represents a distinct chemical space. Performance degradation on non-medicinal datasets, such as ApisTox, demonstrates their limited generalizability of current state-of-the-art algorithms trained solely on biomedical data. Our study highlights the need for more diverse datasets and for targeted model development geared toward the agrochemical domain.
@@
machine learning , chemoinformatics , molecular property prediction , graph neural networks , molecular fingerprints , ecotoxicology , agrochemistry
@@
journal: Computers in Biology and Medicine
@@ 1 Introduction
Low molecular weight compounds, often called small molecules, compounds, or just chemicals, play an important role in various research fields, including material science, biomedical research, and agrochemistry among others. Each field has distinct requirements for such molecules that can help define the chance of their success. These features are often related to physicochemical or molecular properties, for example, in biomedical research the Lipinski Rule of 5 is frequently used to assess lead and drug likeness [1], in agrochemistry similar physicochemical approaches were proposed for different types of pesticides [2, 3] and other types of crop protection compounds[4]. Similarly, important restraints have been established for compound toxicity, allowing both prediction [5, 6] and legislation considering harmful substances (e.g. REACH).
@@
The agrochemistry field shares similarities with biomedical research, often connected with evolutionary conserved mechanisms of actions or even molecular targets. However, significant differences between plants, insects, fungi, animals, and humans are frequently exploited when designing selective and specific small chemicals and biologics. Despite the abundance of medicinal chemistry datasets and benchmarks, such as MoleculeNet [7], PDBBind [8] and Therapeutics Data Commons (TDC) [9], curated agrochemical and environmental datasets remain scarce. This presents a significant challenge for developing robust predictive models. For example, while there are numerous models of human toxicity capable of predicting the off-target effects of known chemicals, there are few analogous models for agrochemicals [10]. This gap is especially pronounced in the prediction of toxicity for environmentally important insects such as bees and specific water organisms, or economically important plants like crops.
@@
Large-scale worldwide use of herbicides and other pesticides can cause toxicities to non-target organisms and environmental degradation. A growing number of studies highlight the need for more selective herbicides targeting plant-specific pathways without affecting aquatic species, microbiota, insects, or mammals. [11][12][13] Research on agrochemical toxicity has become even more essential with respect to environmentally significant insects like honey bee (Apis mellifera), which greatly influence both agriculture health and economies. This important pollinator, vulnerable to various chemicals, highlights both the necessity for more precise and species-specific toxicity data, and for more robust predictive models to assess and mitigate toxicity risks.
@@
The main scope of this research is the exploration of ApisTox, the most comprehensive dataset on bee toxicity [14] with various machine learning approaches to establish current predictive capabilities in this important domain. ApisTox includes chemical compounds represented as SMILES strings, along with their experimentally determined toxicity data for Apis mellifera. We assess the potential of ApisTox for the development of predictive methods ranging from molecular fingerprints to graph kernels, to graph neural networks, and pretrained neural models. Next, we analyze this dataset in terms of its uniqueness and similarities with respect to selected medicinal datasets included in the MoleculeNet benchmark. Finally, we draw conclusions on what works, what fails, and what types of deficiencies need to be addressed in order for the field of bee toxicity prediction to move forward. We explore the limits of current state-of-the-art machine learning (ML) methods, including comparisons with established methods such as molecular fingerprints or graph kernels. Explainable AI tools are also employed for behavioral testing, to verify the chemical soundness of predictions. s such, this work provides value not only to the ML community but also to experimental researchers and entities involved in toxicity studies, both from scientific and applied perspectives. Our study aims to contribute to ongoing efforts to reduce animal testing burdens and to support the development of safer agrochemicals, protect essential pollinators like bees, and promote sustainable agricultural practices.
@@ 2.1 ApisTox dataset
We focus on a recently published ApisTox dataset [14], which provides toxicity data on pesticides and other agrochemicals for the honey bee (Apis Mellifera). ApisTox consists of 1035 compounds, taken from the ECOTOX, PPDB, and BPDB databases, with an applied deduplication and standardization pipeline. This makes it larger and considerably higher quality than previous ML datasets in this area, such as BeeTOX [15]. Notably, it contains a much higher number of 296 toxic molecules, as well as 739 non-toxic ones. The dataset is therefore moderately imbalanced, with 29% of pesticides being toxic. It is overall quite large, compared to most targeted agrochemical datasets and contains a very significant proportion of currently used pesticides. As such, it is highly representative of the overall pesticide chemical space while also providing annotations for the bee toxicity target.
Uniquely, ApisTox provides predetermined train-test splits, which ensures a fair comparison of different algorithms: maximum diversity split (MaxMin) [14, 16], as well as approximation of time split, based on PubChem literature reference. Those split approaches are shown to be potentially quite challenging for ML algorithms, with diverse test sets and high train-test separation, avoiding data leakage related to molecular similarity. This allows us to compare various ML-based molecular property prediction methods and extend this benchmark in the future.
We focus on the binary classification task of distinguishing toxic vs non-toxic molecules. The labels in ApisTox denote the most serious toxicity type, i.e. contact, oral, or other. As such, molecules in the positive class are highly toxic in at least one way to honey bees and can be considered unsafe to use.
@@ 2.2 Molecular dataset analysis methods
The original ApisTox publication [14] contains many analyzes, in particular showing that toxic and non-toxic pesticides vary in terms of molecular structure but not in basic physicochemical properties. In this work, we further analyze the properties of this dataset, showing that it covers a specific agrochemistry-related chemical space. This sets it apart from most benchmarks in molecular property prediction, which are focused on medicinal chemistry, e.g. MoleculeNet [7] and Therapeutics Data Commons (TDC) [9]. In this section, we describe the chemoinformatics tools used for analysis of ApisTox and inter-dataset differences.
@@ 2.2.1 Molecular filters
Molecular filters are a set of rules used to select and limit a set of chemical molecules, which leads to obtaining the desired chemical space. They are used, among other things, in molecular screening to reduce the number of compounds to be analyzed by eliminating undesirable structures, such as those with low bioavailability or potential toxicity. Molecular filters can be divided into two main categories: structural filters, which classify molecules on the basis of the presence of specific functional groups or chemical fragments, and physicochemical filters, which evaluate compounds for selected properties such as molecular weight, logP, or the number of hydrogen bond donors and acceptors. A molecule passes the filter if it meets its criteria, e.g. does not contain any problematic groups, or has the properties in desired ranges. Molecular filter analysis allows the characterization and approximation of the types of chemical structures present in a given set of molecules [17].
@@ 2.2.2 Molecular diversity
One of the simplest methods of analyzing such properties is comparing the distribution or types of chemical elements present in molecules, as well as substructures, functional groups, or fragments. A slight problem with such an analysis is that there is no single commonly accepted definition of a functional group or chemical fragment in the literature. Different approaches are usually used, based on different sets of functional groups, typically defined and published as collections of SMARTS patterns. Such methods are also often used as molecular fingerprints for building QSAR predictive models. Examples include the Laggner fingerprint [20] and MACCS Keys [21], but there are also more sophisticated approaches, like Ertl’s algorithm [22].
Arguably, the most popular method of comparing chemical spaces of datasets is checking the average pairwise Tanimoto of similarity between their molecules. Molecular fingerprints are typically used as a vectorization method, e.g. ECFP4 hashed fingerprint [23], or substructural PubChem fingerprint [24]. The Tanimoto coefficient is considered the most suitable similarity measure for fingerprint-based molecular comparison [25].
ECFP4 fingerprint [23] is most commonly used as a vectorization method. It is an intuitive measure, with interpretable values. However, this approach has some downsides, such as being susceptible to clumping (clustering) of molecules [26], and is often only modestly correlated with biological activity [27].
#Circles measure [26, 27] has been proposed as a solution to these problems. Inspired by sphere exclusion clustering , it measures the coverage of the chemical space by identifying the local neighborhoods occupied by a given set of molecules. It counts the maximum number of mutually exclusive circles (hyperspheres) with a given radius in a dataset. Formally, given a set $\mathcal{S}$ with $n$ molecules, distance measure $d$ and threshold $t\in\left[0,1\right)$, we define #Circles measure as:
#$\#Circles(\mathcal{S},d,t)=\max_{\mathcal{C}\subseteq\mathcal{S}}\left|\mathcal{C}\right|\quad\text{where}\quad d(x,y)\geq t\quad\forall x\neq y\in\mathcal{C}$ (1)
@@ 2.3.1 Molecular fingerprints
Substructure fingerprints extract a predefined set of subgraphs, e.g. functional groups, aromatic rings, or atoms of a given element. The choice of substructures is typically selected by domain experts, most typically medicinal chemists. One can check the existence of a given substructure (binary fingerprint) or count its occurrences in a molecule (count fingerprint). Examples include MACCS Keys [21] and PubChem fingerprint [24]. They are very interpretable and can encode domain-relevant features.
Hashed fingerprints define the general shape of subgraphs that are extracted from a molecule. Those can be, for example, circular neighborhoods for the ECFP fingerprint [23] or shortest paths between pairs of atoms in the Atom Pair fingerprint [29]. A unique identifier is assigned to each subgraph, which is then hashed onto a constant-length output vector at a position based on the identifier value. In binary fingerprints, we ignore hashing collisions, whereas count fingerprints sum up all occurrences at a given index. They are the most flexible group, often resulting in very strong ML classifiers, competitive with complex GNNs [30, 31].
@@ 2.3.3 Graph kernels
Graph kernels quantify the similarity between graphs, based on their structure [35, 36, 37]. In many cases, they also allow for incorporating vertex attributes, such as element type, and many molecule kernels have been proposed to describe the structural similarity of compounds. The resulting kernel matrix, containing pairwise similarity values between molecules, allows for the use of any established kernel methods, such as the SVM classifier. Due to their high expressive power, they show excellent performance in chemoinformatics [38, 39, 40].
The simplest possible kernels are the vertex and edge histogram kernels [41], which use a dot product between the counts of elements and the bonds between compounds, respectively. Many more complex kernels have been proposed, capturing local and global topology of graphs, e.g. shortest paths [42] or graphlets (all subgraphs up to a given size) [43, 44]. The propagation kernel [45] has been proposed to diffuse the node information in the graph, strongly focusing on the functional similarity between molecules, in addition to their topology.
The Weisfeiler-Lehman (WL) kernel [46] is very commonly used, due to its excellent performance, speed, and strong theoretical properties. It is based on the Weisfeiler-Lehman isomorphism test [47, 48], aimed at distinguishing non-isomorphic graphs. The idea is to exchange information between neighboring nodes (bonded atoms), which captures the topological structure of a $k$-hop neighborhood around each atom after $k$ iterations. In particular, it has also directly influenced the Morgan algorithm for the molecular graph isomorphism test [49] and the ECFP fingerprint [23]. This isomorphism testing framework also underlies the theory behind message-passing GNNs. The WL kernel is particularly computationally efficient, compared to most graph kernels.
The Weisfeiler-Lehman Optimal Assignment (WL-OA) kernel [50] has been proposed as an extension. The idea is that two similar molecules should have matching parts, e.g. same or similar rings or functional groups. This is realized by the optimal assignment kernel [51], which acts as a function mapping atoms in two molecules (identified using the Weisfeiler-Lehman procedure) to maximize the sum of the pairwise atom similarities. This substantially strengthens the ability of the kernel to distinguish between similar and dissimilar molecular substructures. Thus, it achieves very strong classification performance, albeit at a considerably higher computational cost.
@@ 2.3.4 Graph neural networks (GNNs)
Graph neural networks (GNNs) [52, 53, 54, 55] are a family of neural networks for graph representation learning, with appropriate inductive biases and theoretical guarantees. In particular, their operations, like graph convolution, are permutation-invariant, allowing them to process vertices and edges without assuming any natural order. They also naturally incorporate any atom and bond features. This makes them a natural choice for learning embeddings of molecules.
@@
$k$-layer GNN is able to aggregate information about the $k$-hop neighborhood around each atom in its embedding. This is important for molecular classification since, e.g. understanding ring structures requires many layers under this paradigm. However, building deep GNNs has proven difficult, due to oversmoothing, oversquashing, and unstable training [61], which make it difficult to capture long-range dependencies. Jumping Knowledge [62] has been proposed as a simple solution, where we add skip connections, utilizing atom embeddings after each layer to obtain the molecule embedding, instead of just vectors after the final layer. Normalization layers, such as LayerNorm [63, 64, 65] can help stabilize training. Bayesian hyperparameter optimization (HPO) [66], e.g. using Parzen tree estimator [67], can be used to efficiently select good hyperparameter values, like learning rate or number of layers, enhancing the performance of the resulting architecture.
AttentiveFP [66] applied many of these ideas specifically for molecular property prediction. It utilizes the attention mechanism combined with artificial bonds for capturing long-range dependencies. It also uses the GRU recurrent network [68] for graph level readout, to focus on the most relevant atom embeddings. They also utilized Bayesian HPO to tune many hyperparameters of this architecture.

### v4

# Evaluating machine learning models for predicting pesticide toxicity to honey bees

@@ Abstract
Small molecules play a critical role in the biomedical, environmental, and agrochemical domains, each with distinct physicochemical requirements and success criteria. Although biomedical research benefits from extensive datasets and established benchmarks, agrochemical data remain scarce, particularly with respect to species-specific toxicity. This work focuses on ApisTox, the most comprehensive dataset of experimentally validated chemical toxicity to the honey bee (Apis mellifera), an ecologically vital pollinator. The primary goal of this study was to determine the suitability of diverse machine learning approaches for modeling such toxicity, including molecular fingerprints, graph kernels, and graph neural networks, as well as pretrained models. Comparative analysis with medicinal datasets from the MoleculeNet benchmark reveals that ApisTox represents a distinct chemical space. Performance degradation on non-medicinal datasets, such as ApisTox, demonstrates their limited generalizability of current state-of-the-art algorithms trained solely on biomedical data. Our study highlights the need for more diverse datasets and for targeted model development geared toward the agrochemical domain.
@@
agrochemistry , chemoinformatics , machine learning , molecular fingerprints , molecular property prediction
@@
journal: Ecotoxicology and Environmental Safety
@@ 1 Introduction
Low molecular weight compounds, often called small molecules, compounds, or just chemicals, play an important role in various research fields, including material science, biomedical research, and agrochemistry among others. Each field has distinct requirements for such molecules that can help define the chance of their success [7, 71]. These features are often related to physicochemical or molecular properties, for example, in biomedical research the Lipinski Rule of 5 is frequently used to assess lead and drug likeness [31], in agrochemistry similar physicochemical approaches were proposed for different types of pesticides [21, 56] and other types of crop protection compounds [72]. Similarly, important restraints have been established for compound toxicity, allowing both prediction [14, 6] and legislation considering harmful substances (e.g. REACH).
@@
The agrochemistry field shares similarities with biomedical research, often connected with evolutionary conserved mechanisms of actions or even molecular targets [62, 53]. However, significant differences between plants, insects, fungi, animals, and humans are frequently exploited when designing selective and specific small chemicals and biologics [57, 40]. Despite the abundance of medicinal chemistry datasets and benchmarks, such as MoleculeNet [64] and Therapeutics Data Commons (TDC) [24], curated agrochemical and environmental datasets and benchmarks, ready to use for machine learning algorithms, remain scarce. Large-scale resources like ECOTOX [41] require significant data curation efforts, such as structural standardization, deduplication, and merging multiple experimental measurements for each compound into a single label for supervised learning [4]. This presents a significant challenge for developing robust predictive models. For example, while there are numerous models of human toxicity capable of predicting the off-target effects of known chemicals, there are few analogous models for agrochemicals [16]. This gap is especially pronounced in the prediction of toxicity for environmentally important insects such as bees and specific water organisms, or economically important plants like crops.
@@
Large-scale worldwide use of herbicides and other pesticides can cause toxicities to non-target organisms and environmental degradation. A growing number of studies highlight the need for more selective herbicides targeting plant-specific pathways without affecting aquatic species, microbiota, insects, or mammals. [32, 22, 13] Research on agrochemical toxicity has become even more essential with respect to environmentally significant insects like honey bee (Apis mellifera), which greatly influence both agriculture health and economies. This important pollinator, vulnerable to various chemicals, highlights both the necessity for more precise and species-specific toxicity data, and for more robust predictive models to assess and mitigate toxicity risks.
@@
The main scope of this research is the exploration of ApisTox, the most comprehensive dataset on bee toxicity [4] with various machine learning approaches to establish current predictive capabilities in this important domain. ApisTox includes chemical compounds represented as SMILES strings, along with their experimentally determined toxicity data for Apis mellifera. The potential of ApisTox for the development of predictive methods is assessed, ranging from molecular fingerprints to graph kernels, to graph neural networks, and pretrained neural models. Next, we analyze this dataset in terms of its uniqueness and similarities with respect to selected medicinal datasets included in the MoleculeNet benchmark. Finally, we draw conclusions on what works, what fails, and what types of deficiencies need to be addressed in order for the field of bee toxicity prediction to move forward. Limits of current state-of-the-art machine learning (ML) methods are explored, including comparisons with established methods such as molecular fingerprints or graph kernels. Explainable AI tools are also employed for behavioral testing, to verify the chemical soundness of predictions. As such, this work provides value not only to the ML community but also to experimental researchers and entities involved in toxicity studies, both from scientific and applied perspectives. This study aims to contribute to ongoing efforts to reduce animal testing burdens and to support the development of safer agrochemicals, protect essential pollinators like bees, and promote sustainable agricultural practices.
@@ 2.1 ApisTox dataset
The main data source used in this study is a recently published ApisTox dataset [4], which provides toxicity data on pesticides and other agrochemicals for the honey bee (Apis Mellifera). ApisTox consists of 1035 compounds, taken from the ECOTOX, PPDB, and BPDB databases, with an applied deduplication and standardization pipeline. This makes it larger and considerably higher quality than previous ML datasets in this area. Notably, it contains a much higher number of 296 toxic molecules, as well as 739 non-toxic ones. The dataset is therefore moderately imbalanced, with 29% of pesticides being toxic. It is overall quite large, compared to most targeted agrochemical datasets in this area prepared for ML modeling [4], e.g. CropCSM [45] (871 molecules), BeeTOX [61] (876 molecules), or BeeToxAI [36] (472 molecules). Further, ApisTox contains a very significant proportion of currently used pesticides. As such, it is highly representative of the overall pesticide chemical space while also providing annotations for the bee toxicity target.
Uniquely, ApisTox provides predetermined train-test splits, which ensures a fair comparison of different algorithms: maximum diversity split (MaxMin) [4], as well as approximation of time split, based on PubChem literature reference. Those split approaches are shown to be potentially quite challenging for ML algorithms, with diverse test sets and high train-test separation, avoiding data leakage related to molecular similarity. This allows comparing various ML-based molecular property prediction methods and extend this benchmark in the future.
In machine learning terms, the task in this work is binary classification of toxic vs non-toxic molecules, primarily based on their molecular graph representation. The labels in ApisTox denote the most serious toxicity type, i.e. contact, oral, or other. As such, molecules in the positive class are highly toxic in at least one way to honey bees and can be considered unsafe to use.
@@ 2.2 Molecular dataset analysis methods
The original ApisTox publication [4] contains many analyses, in particular showing that toxic and non-toxic pesticides vary in terms of molecular structure but not in basic physicochemical properties. In this work, we further analyze the properties of this dataset, showing that it covers a specific agrochemistry-related chemical space [62, 71]. This sets it apart from most benchmarks in molecular property prediction, which are focused on medicinal chemistry, e.g. MoleculeNet [64] and Therapeutics Data Commons (TDC) [24]. In this section, we describe the chemoinformatics tools used for analysis of ApisTox and inter-dataset differences.
@@ 2.2.1 Molecular filters
Molecular filters are a set of rules used to select and limit a set of chemical molecules, which leads to obtaining the desired chemical space [27]. They are used, among other things, in molecular screening to reduce the number of compounds to be analyzed by eliminating undesirable structures, such as those with low bioavailability or potential toxicity [27, 60]. Molecular filters can be divided into two main categories: structural filters, which classify molecules on the basis of the presence of specific functional groups or chemical fragments, and physicochemical filters, which evaluate compounds for selected properties such as molecular weight, logP, or the number of hydrogen bond donors and acceptors. A molecule passes the filter if it meets its criteria, e.g. does not contain any problematic groups, or has the properties in desired ranges. Molecular filter analysis allows the characterization and approximation of the types of chemical structures present in a given set of molecules [27].
@@ 2.2.2 Molecular diversity
One of the simplest methods of analyzing such properties is comparing the distribution or types of chemical elements present in molecules, as well as substructures, functional groups, or fragments. A slight problem with such an analysis is that there is no single commonly accepted definition of a functional group or chemical fragment in the literature. Different approaches are usually used, based on different sets of functional groups, typically defined and published as collections of SMARTS patterns. Such methods are also often used as molecular fingerprints for building QSAR predictive models. Examples include the Laggner fingerprint and MACCS Keys.
Arguably, the most popular method of comparing chemical spaces of datasets is checking the average pairwise Tanimoto of similarity between their molecules. Molecular fingerprints are typically used as a vectorization method, e.g. ECFP4 hashed fingerprint, or substructural PubChem fingerprint. The Tanimoto coefficient is the most widely used similarity measure for comparison of resulting fingerprint vectors.
ECFP4 fingerprint [50] is most commonly used as a vectorization method. It is an intuitive measure, with interpretable values. However, this approach has some downsides, such as being susceptible to clumping (clustering) of molecules, and is often only modestly correlated with biological activity [66].
#Circles measure [66] has been proposed as a solution to these problems. Inspired by sphere exclusion clustering, it measures the coverage of the chemical space by identifying the local neighborhoods occupied by a given set of molecules. It counts the maximum number of mutually exclusive circles (hyperspheres) with a given radius in a dataset. Formally, given a set $\mathcal{S}$ with $n$ molecules, distance measure $d$ and threshold $t\in\left[0,1\right)$, we define #Circles measure as:
$Circles(\mathcal{S},d,t)=\max_{\mathcal{C}\subseteq\mathcal{S}}\left|\mathcal{C}\right|\quad\text{where}\quad d(x,y)\geq t\quad\forall x\neq y\in\mathcal{C}$ (1)
@@ 2.3.1 Molecular fingerprints
Substructure fingerprints extract a predefined set of subgraphs, e.g. functional groups, aromatic rings, or atoms of a given element. The choice of substructures is typically selected by domain experts, most typically medicinal chemists. One can check the existence of a given substructure (binary fingerprint) or count its occurrences in a molecule (count fingerprint). Examples include MACCS Keys and PubChem fingerprint. They are very interpretable and can encode domain-relevant features.
Hashed fingerprints define the general shape of subgraphs that are extracted from a molecule. Those can be, for example, circular neighborhoods for the ECFP fingerprint or shortest paths between pairs of atoms in the Atom Pair fingerprint . A unique identifier is assigned to each subgraph, which is then hashed onto a constant-length output vector at a position based on the identifier value. In binary fingerprints, we ignore hashing collisions, whereas count fingerprints sum up all occurrences at a given index. They are the most flexible group, often resulting in very strong ML classifiers, competitive with complex GNNs [3].
@@ 2.3.3 Graph kernels
Graph kernels quantify the similarity between graphs, based on their structure [59]. In many cases, they also allow for incorporating vertex attributes, such as element type, and many molecule kernels have been proposed to describe the structural similarity of compounds. The resulting kernel matrix, containing pairwise similarity values between molecules, allows for the use of any established kernel methods, such as the SVM classifier. Due to their high expressive power, they show excellent performance in chemoinformatics [48].
The simplest possible kernels are the vertex and edge histogram kernels, which use a dot product between the counts of elements and the bonds between compounds, respectively. Many more complex kernels have been proposed, capturing local and global topology of graphs, e.g. shortest paths [9] or graphlets (all subgraphs up to a given size) [47]. The propagation kernel [38] has been proposed to diffuse the node information in the graph, strongly focusing on the functional similarity between molecules, in addition to their topology.
The Weisfeiler-Lehman (WL) kernel [54] is very commonly used, due to its excellent performance, speed, and strong theoretical properties. It is based on the Weisfeiler-Lehman isomorphism test [30], aimed at distinguishing non-isomorphic graphs. The idea is to exchange information between neighboring nodes (bonded atoms), which captures the topological structure of a $k$-hop neighborhood around each atom after $k$ iterations. In particular, it has also directly influenced the Morgan algorithm for the molecular graph isomorphism test and the ECFP fingerprint. This isomorphism testing framework also underlies the theory behind message-passing GNNs. The WL kernel is particularly computationally efficient, compared to most graph kernels.
The Weisfeiler-Lehman Optimal Assignment (WL-OA) kernel [28] has been proposed as an extension. The idea is that two similar molecules should have matching parts, e.g. same or similar rings or functional groups. This is realized by the optimal assignment kernel, which acts as a function mapping atoms in two molecules (identified using the Weisfeiler-Lehman procedure) to maximize the sum of the pairwise atom similarities. This substantially strengthens the ability of the kernel to distinguish between similar and dissimilar molecular substructures. Thus, it achieves very strong classification performance, albeit at a considerably higher computational cost.
@@ 2.3.4 Graph neural networks (GNNs)
Graph neural networks (GNNs) [65] are a family of neural networks for graph representation learning, with appropriate inductive biases and theoretical guarantees. In particular, their operations, like graph convolution, are permutation-invariant, allowing them to process vertices and edges without assuming any natural order. They also naturally incorporate any atom and bond features. This makes them a natural choice for learning embeddings of molecules.
@@
$k$-layer GNN is able to aggregate information about the $k$-hop neighborhood around each atom in its embedding. This is important for molecular classification since, e.g. understanding ring structures requires many layers under this paradigm. However, building deep GNNs has proven difficult, due to oversmoothing, oversquashing, and unstable training [5], which make it difficult to capture long-range dependencies. Jumping Knowledge [69] has been proposed as a simple solution, where we add skip connections, utilizing atom embeddings after each layer to obtain the molecule embedding, instead of just vectors after the final layer. Normalization layers, such as LayerNorm can help stabilize training [33]. Bayesian hyperparameter optimization (HPO), e.g. using Parzen tree estimator [8], can be used to efficiently select good hyperparameter values, like learning rate or number of layers, enhancing the performance of the resulting architecture [67].
AttentiveFP [67] applied many of these ideas specifically for molecular property prediction. It utilizes the attention mechanism combined with artificial bonds for capturing long-range dependencies. It also uses the GRU recurrent network for graph level readout, to focus on the most relevant atom embeddings. They also utilized Bayesian HPO to tune many hyperparameters of this architecture.